OpenStudy (anonymous):

For school, Georgia bought 6 notebooks and 2 folders. Let n represent the cost of each notebook and f represent the cost of each folder. Write an expression to determine the total cost for the school supplies. Then, calculate the cost if each notebook costs fifty cents and each folder costs two dollars.

OpenStudy (anonymous):

Through a bit of algebra, we can write this expression \[0.5n + 3f\] we have 6 notebooks (n = 6) and 2 folders (f = 2) in dollars, we get 0.5(6) + 3(2) = 3 + 6 = 9.00 in dollars

OpenStudy (tkhunny):

Homw many notebooks and folders did South Carolina buy? That just doesn't seem like enough for a whole State!

OpenStudy (anonymous):

theses are the answers though two n plus six f; seven dollars two n plus six f; thirteen dollars six n plus two f; thirteen dollars six n plus two f; seven dollars

OpenStudy (anonymous):

@hfw